# Microquasar jet-cocoon systems as PeVatrons

### v1

@@ III Observed cosmic-ray spectrum on Earth
Figure 3: Predicted CR intensity on Earth and the mean mass $\langle{\rm ln}(A)\rangle$ distribution from microquasars. The all-particle energy spectra are taken from LHAASO Cao et al. (2024a), HAWC Alfaro et al. (2017), NUCLEON Grebenyuk et al. (2019), TIBET Amenomori et al. (2008), TA Matthews (2017), and Auger Verzi (2019), while the proton spectrum are taken from LHAASO Cao et al. (2025), KASCADE-Grande Arteaga-Velázquez et al. (2017), and IceTop Aartsen et al. (2019). The measurement of the mean mass $\langle{\rm ln}(A)\rangle$ distribution are taken from LHAASO Cao et al. (2024a) and IceTop Aartsen et al. (2019).
@@
[…]
@@
where $N_{i}(E,\mathbf{x})$ is the number density per energy at position $\mathbf{x}$, the second and third term describes the time evolution of particle number density due to advection and spatial diffusion, respectively. The first term in the right-hand represents injection of particles of type $i$, and the second and last term accounts for the loss and gain processes. Here, the loss of nuclei of type $i$ caused by collisions and decay with a rate $p_{i}=v\rho/\lambda_{i}+1/\gamma\tau_{i}$, where $\rho$ is the interstellar medium density, $\lambda_{i}$ is the mean free path of interactions when particle propagates inside Galaxy, $\tau_{i}$ is the decay time of particle of type $i$, and $\gamma$ is particle Lorentz factor. In this study, we employ a publicly available version of the DRAGON code to propagate CRs from sources to Earth Evoli et al. (2017). The maximum value of Galactocentric radius is $R_{G}=10\rm~{}kpc$ and halo size is $H_{G}=4\rm~{}kpc$. The spatial distribution of microquasars is unknown. For microquasars like SS 433 or Sgr V4641, it follows the distribution of massive stars, thus have similar spatial distribution with Galactic SNRs. For the source distribution, we adopt the Ferriere2001 model Ferriere (2001), based on the population studies on SNRs. The spatial distribution of the diffusion coefficient is constant and parametrized as $D=23.8\times 10^{28}(\mathcal{R}/250\rm~{}GV)^{\delta}{\rm cm}^{2}{\rm~{}s}^{-1}$ with $\delta=0.45$ for $\mathcal{R}<250\rm~{}GV$, while $\delta=1/3$ for $\mathcal{R}\geqslant 250\rm~{}GV$ Murase and Fukugita (2019); Kimura et al. (2018b). The chosen diffusion coefficient is determined by the observed Boron to Carbon (B/C) ratio.
@@
We adjust the energy injection rate of CRs to match the CR data at PeV energies, which requires $Q_{\rm cr,tot}\sim 1.3\times 10^{42}\rm~{}erg~{}kpc^{-3}~{}yr^{-1}$. The energy injection rate integrated over the Galactic volume is $L_{\rm cr,tot}=Q_{\rm CR,tot}V_{G}\simeq 1.5\times 10^{38}\rm~{}erg~{}s^{-1}$, where $V_{G}\sim 2\pi R_{G}^{2}\times H_{G}\sim 1\times 10^{68}\rm~{}cm^{3}$ is the Galactic volume. Our scenario is consistent with the value required to explain CRs in the knee energy range if we have around $3-10$ microquasars in our Galaxy. For our scenario to work, microquasars need to have extended jets of $\sim 100$ pc. Only two objects (SS433 and V4641 Sgr) exhibit such a feature. The other 3 microquasars, GRS 1915, Cyg X-1, MAXI J1820, do not have such large-scale extended jets, although they have more compact jets. Since we can observe only less than half of our Galaxy, 3-10 microquasars with extended jets sounds more reasonable. Note that the required number of microquasars with extended jets depend on the injection energy given by Eq. II.2.
@@
In Fig. 3, we present the predicted energy spectrum and mean mass $\langle{\rm ln}(A)\rangle$ of CRs from Galactic microquasars in our shear acceleration scenario. Note the shaded region in the lower panel about $\langle{\rm ln}(A)\rangle$ are affected by low- and high-energy components. In our scenario, the CR composition ratio is determined by the Galactic CR composition ratio at the rigidity of $\mathcal{R}_{\rm inj}=E_{\rm inj}/Z$. Since $E_{\rm inj}$ is uncertain, we examine several values of $E_{\rm inj}$ (see the next paragraph). In Fig. 3, we adopt the CR composition at 300 GV from AMS-02 observation  Aguilar et al. (2018a, b, 2015, 2021, 2020). Our model demonstrates that CRs from microquasars could explain both proton and all-particle spectrum around knee. The hard proton spectrum below the peak energy (at a few PeV) and a steep power-law decay beyond the peak predicted by our model agrees well with the LHAASO observations. Furthermore, the model self-consistently predicts the spectra of helium, intermediate-mass, and heavy nuclei around knee, with their relative abundance normalized to the local Galactic CR abundance at 100 GV. Future precision measurements of light and heavy nuclei will provide a critical test of this scenario.
@@
For comparison, we also adopt the CR composition inferred from AMS-02 observations at reference rigidities of 100 GV and 1 TV Aguilar et al. (2018a, b, 2015, 2021, 2020). We also consider a scenario where injected CRs follow solar abundances, $f_{\odot}$, as expected if they are swept from the interstellar medium (ISM) Lodders (2010). The adopted composition ratios are summarized in Table 2. As demonstrated in Ref. Caprioli et al. (2017), the nonthermal tail of nuclei is enhanced by a factor of $(A/Z)^{2}$ for efficient diffusive acceleration in nonrelativistic shocks. Our analysis shows that reaccelerating low-energy CRs (100 GV – 1 TV) yields a mean logarithmic mass $\langle\ln A\rangle$ roughly consistent with observations near the CR knee. In contrast, models assuming solar or enhanced solar compositions fail to reproduce the observed mass distribution. These results might indicate that the reacceleration of pre-existing low-energy CRs provides a plausible explanation for the composition trends observed in the knee region.
Z A $f_{A,\rm 300GV}$ $f_{A,\rm 100GV}$ $f_{A,\rm 1TV}$ $f_{\odot}$ $f_{\odot}^{\rm enhance}$
@@
Table 2: Injected CR composition at the rigidity 300 GV, 100 GV, and 1 TV for different composition models. In addition, we also include solar composition $f_{\odot}$ and enhanced solar composition $f_{\odot}^{\rm enhance}$ by a factor of $(A/Z)^{2}$.
@@ IV Connection to low- and high-energy components
In this section, we aim to demonstrate that our microquasar model can be combined with other astrophysical models explaining the CR spectrum at different wavelengths. In particular, we provide a possible model, in which the whole CR spectrum from around GeV to 100 EeV can be explained with four main sources, including supernova remnants (SNR, Gal pop1), microquasar (MQ, Gal pop2), clusters of galaxies (CG, ExGal pop1) and radio galaxies (RQ, ExGal pop2). Note that the main purpose here is to demonstrate that the reacceleration model works not only in the large-scale jet-cocoon system of microquasars but also in radio galaxies.
@@
For CRs from Galactic SNRs, we calculate the observed energy spectrum of each element with DRAGON code, and the normalization is adjusted to match the AMS-02 data. We adopt an broken-power law injection spectrum of CRs from SNRs to fit the data. See Appendix for the details. The escaped CR spectrum typically have exponential cutoff beyond the maximum acceleration energy. We set the maximum rigidity from Galactic SNRs to $\mathcal{R}_{\rm max}=0.4\rm~{}PV$.
@@
Figure 4: An overall picture of the CR spectrum from Galactic to extragalactic components. Gal pop1: CRs from supernova remnants (SNR) Reynolds (2008); Blasi (2013) (dotted curve); Gal pop2: CRs from microquasars (MQ) (dashed curve); ExGal pop1: CRs from extragalactic sources, e.g., cluster of galaxies Murase et al. (2008) (dotted-dashed curve); ExGal pop2: UHECRs from radio galaxies (RG) Kimura et al. (2018a) (long dashed curve). The data points are the same as those in Fig. 3, with the AMS data are taken from Refs. Aguilar et al. (2018a, b, 2015, 2021, 2020), and the mean mass distribution measured by Auger is taken from Ref. Yushkov (2020).
@@
For CRs in the knee energy range, we use the CR spectrum that comes directly from Monte Carlo simulations, same as Fig. 3. At energies above around 100 PeV to EeV, CRs from extragalactic sources begin to dominate. At this energy range, the second knee appears, where the spectrum is further steepened from $\gamma\sim 3$ to $\gamma\sim 3.3$, because of the limitation of the maximum energy of the second Galactic component based on the rigidity-dependent modeling. The physical origin of the CRs beyond the second knee up to the ankle has been proposed by various models, and it has been shown that an additional component so-called the B-component is required Gaisser et al. (2016). This component may be responsible for the proton subankle (hardening) feature around 100 PeV observed by KASCADE-Grande Apel et al. (2013) , and the proposed models include the cluster of galaxies Murase et al. (2008, 2013); Simeon et al. (2025), Galactic wind Thoudam et al. (2016); Mukhopadhyay et al. (2023), Wolf-Rayet star explosions Thoudam et al. (2016). In addition, super-knee CR accelerators such as Galactic neutron star merger remnants Kimura et al. (2018b) and interacting supernovae Ptuskin and Zirakashvili (2003, 2005) may explain another hardening feature around $10^{16.22}$ eV reported by TALE Abbasi et al. (2018).
@@
For the subankle proton component, we consider cluster of galaxies for the demonstrative purpose. We consider the injected CRs have composition similar to solar composition, but enhanced by a factor of $(A/Z)^{2}$, see the last column in Tabel 2 about $f_{\odot}^{\rm enhance}$. Note that our results are unaffected by the composition assumption unless the CRs are dominated by heavy nuclei, and such nuclei can be largely disintegrated inside the clusters Kotera et al. (2009); Fang and Murase (2018). The source spectral index above 10-100 PeV is $s_{\rm cr}=2.4$ and the maximum acceleration rigidity $\mathcal{R}_{\rm max}=3.7\times 10^{17}\rm~{}eV$, following Ref. Murase et al. (2008). We propagate CRs from an ensemble of extragalactic sources that are uniformly distributed with the characteristic distance between sources $d_{s}=32.5\rm~{}Mpc$. The number density of galaxy clusters is around $10^{-6}\rm~{}Mpc^{-3}$, which leads to $d_{s}\sim 100$ Mpc. The observed flux is obtained with Eq. 6 in Ref. González et al. (2021). We assume a uniform extragalactic magnetic field $B_{\rm rms}=1\rm~{}nG$ and coherence length $L_{\rm coh}=1\rm~{}Mpc$.
@@
UHECRs above EeV energies are produced from radio galaxies via the shear-espresso acceleration mechanism Caprioli (2015); Kimura et al. (2018a); Mbarek and Caprioli (2019); Mbarek et al. (2023); Seo et al. (2024); Wang et al. (2024); Mbarek et al. (2025). Here, we adopt the predicted spectrum and composition model of UHECRs in KMZ18, where CRs with rigidity $\sim 1$ TV in radio galaxies are re-accelerated by shear in kpc-scale jets. Since radio galaxies are more evolved than our Galaxy, KMZ18 considers that the metal composition is enhanced by a ratio of 3, compared to our Galaxy. Here, we further enhance the iron composition ratio by a factor of 3 to improve the fit to the latest UHECR data. The further enhancement of iron composition ratio in radio galaxies is reasonable because elliptical galaxies have lower star-formation rates than our Galaxy. In radio galaxies, core-collapse supernovae do not frequently occur because of their low star-formation rates, while type Ia supernovae occur because explosions of white dwarfs require significant time lag after star-formation activities. Then, type Ia supernovae inject iron into the ISM, enhancing the iron composition ratio. In addition, we consider a larger value of magnetic field strength in the cocoon with $B_{\rm coc}=4.5\rm~{}\mu G$, where the peak energy of the injection spectrum shown in Fig. 3 in KMZ18 is increased to $\mathcal{R}_{\rm max}\approx 2.4\rm~{}EV$
@@
We found that the overall CR spectrum and composition could be explained by the above four populations. However, one should keep in mind that uncertainty is currently large, and this demonstrates that the reacceleration of Galactic CRs plays a key role in reshaping the mass distribution of accelerated CRs at different energy ranges.

### v2

@@ III Observed cosmic-ray spectrum on Earth
Figure 3: Predicted CR intensity on Earth and the mean mass $\langle{\rm ln}(A)\rangle$ distribution from microquasars. The all-particle energy spectra are taken from LHAASO Cao et al. (2024a), HAWC Alfaro et al. (2017), NUCLEON Grebenyuk et al. (2019), TIBET Amenomori et al. (2008), TA Matthews (2017), and Auger Verzi (2019), while the proton spectra are taken from LHAASO Cao et al. (2025), KASCADE-Grande Arteaga-Velázquez et al. (2017), and IceTop Aartsen et al. (2019). The measurements of the mean mass $\langle{\rm ln}(A)\rangle$ distribution are taken from LHAASO Cao et al. (2024a) and IceTop Aartsen et al. (2019).
@@
$+\frac{v\rho(\mathbf{x})}{m_{p}}\sum_{k\geq i}\int\frac{d\sigma_{i,k}(E,E^{\prime})}{dE}N_{k}(E^{\prime},\mathbf{x})dE^{\prime},$ (6)
@@
where $N_{i}(E,\mathbf{x})$ is the number density per energy at position $\mathbf{x}$, and the second and third terms on the left side describe the time evolution of the particle number density due to advection and spatial diffusion, respectively. The first term on the right side represents the injection of particles of type $i$, and the second and last terms account for the loss and gain processes. Here, the loss of nuclei of type $i$ caused by collisions and decay with a rate $p_{i}=v\rho/\lambda_{i}+1/\gamma\tau_{i}$, where $\rho$ is the interstellar medium density, $\lambda_{i}$ is the mean free path of interactions when particle propagates inside the Galaxy, $\tau_{i}$ is the decay time of a particle of type $i$, and $\gamma$ is theparticle Lorentz factor. In this study, we employ a publicly available version of the DRAGON code to propagate CRs from sources to Earth Evoli et al. (2017). The maximum value of the galactocentric radius is $R_{G}=10\rm~kpc$ and the halo size is $H_{G}=4\rm~kpc$. The spatial distribution of microquasars is unknown. High-Eddington ratio microquasars, including SS 433 or Sgr V4641, likely follow the distribution of massive stars, and thus, they have similar spatial distribution with Galactic SNRs. For the source distribution, we adopt the Ferriere2001 model Ferriere (2001), based on the population studies of SNRs. The spatial distribution of the diffusion coefficient is constant and parametrized as $D=23.8\times 10^{28}(\mathcal{R}/250\rm~GV)^{\delta}{\rm cm}^{2}{\rm~s}^{-1}$ with $\delta=0.45$ for $\mathcal{R}<250\rm~GV$, while $\delta=1/3$ for $\mathcal{R}\geqslant 250\rm~GV$ Murase and Fukugita (2019); Kimura et al. (2018b). The chosen diffusion coefficient is determined by the observed boron to carbon (B/C) ratio Aguilar et al. (2016).
We adjust the energy injection rate of CRs to match the CR data at PeV energies, which requires $Q_{\rm cr,tot}\sim 1.3\times 10^{42}\rm~erg~kpc^{-3}~yr^{-1}$. The energy injection rate integrated over the Galactic volume is
$L_{\rm cr,tot}=Q_{\rm CR,tot}V_{G}\simeq 1.5\times 10^{38}\rm~erg~s^{-1},$ (7)
@@
where $V_{G}\sim 2\pi R_{G}^{2}\times H_{G}\sim 1\times 10^{68}\rm~cm^{3}$ is the Galactic volume. Our scenario is consistent with the value required to explain CRs in the knee energy range if we have around three to ten microquasars in our Galaxy. For our scenario to work, microquasars need to have extended jets of $\sim 100$ pc. Only two objects (SS433 and V4641 Sgr) exhibit such a feature. The other three microquasars, GRS 1915, Cyg X-1, MAXI J1820, do not have such large-scale extended jets, although they have more compact jets. Since we can observe only less than half of our Galaxy, three to ten microquasars with extended jets sounds more reasonable. Note that the required number of microquasars with extended jets depends on the injection energy given by Eq. II.2.
@@
In Fig. 3, we present the predicted energy spectrum and mean mass $\langle{\rm ln}(A)\rangle$ of CRs from Galactic microquasars in our shear acceleration scenario. Note the shaded region in the lower panel about $\langle{\rm ln}(A)\rangle$ are affected by low- and high-energy components. In our scenario, the CR composition ratio is determined by the Galactic CR composition ratio at the rigidity of $\mathcal{R}_{\rm inj}=E_{\rm inj}/Z$. Since $E_{\rm inj}$ is uncertain, we examine several values of $E_{\rm inj}$ (see the next paragraph). In Fig. 3, we adopt the CR composition at 300 GV from the AMS-02 observation  Aguilar et al. (2018a, b, 2015, 2021, 2020). Our model demonstrates that CRs from microquasars could explain both proton and all-particle spectrum around the knee. The hard proton spectrum below the peak energy (at a few PeV) and a steep power-law decay beyond the peak predicted by our model agrees well with the LHAASO observations. Furthermore, the model self-consistently predicts the spectra of helium, intermediate mass, and heavy nuclei around knee, with their relative abundance normalized to the local Galactic CR abundance at 300 GV. Future precision measurements of light and heavy nuclei will provide a critical test of this scenario.
@@
For comparison, we also adopt the CR composition inferred fromt AMS-02 observations at reference rigidities of 100 GV and 1 TV Aguilar et al. (2018a, b, 2015, 2021, 2020). We also consider a scenario where injected CRs follow solar abundances, $f_{\odot}$, as expected if they are swept from the interstellar medium (ISM) Lodders (2010). The adopted composition ratios are summarized in Table 2. As demonstrated in Ref. Caprioli et al. (2017), the nonthermal tail of nuclei is enhanced by a factor of $(A/Z)^{2}$, $f_{\odot}^{\rm enhance}$, for efficient diffusive acceleration in nonrelativistic shocks considering fully ionized state. Our analysis shows that reaccelerating low-energy CRs (100 GV – 1 TV) yields a mean logarithmic mass $\langle\ln A\rangle$ roughly consistent with observations near the CR knee. In contrast, models assuming solar or enhanced solar compositions fail to reproduce the observed mass distribution. For CRs from microquasars, the $(A/Z)^{2}\sim A^{2}$ enhancement scenario – expected in a singly ionized medium – predicts a heavier CR composition than a fully ionized medium. However, this leads to an overly heavy composition, inconsistent with the observed $\langle\ln A\rangle$. These results might indicate that the reacceleration of preexisting low-energy CRs provides a plausible explanation for the composition trends observed in the knee region.
Table 2: Injected CR composition at the rigidity 300 GV, 100 GV, and 1 TV for different composition models. In addition, we also include solar composition $f_{\odot}$ and enhanced solar composition $f_{\odot}^{\rm enhance}$ by a factor of $(A/Z)^{2}$ for a fully ionized medium.
Z A $f_{A,\rm 300GV}$ $f_{A,\rm 100GV}$ $_{A,\rm 1TV} ff_{\odot}$ $f_{\odot}^{\rm enhance}$
@@ IV Connection to low- and high-energy components
In this section, we aim to demonstrate that our microquasar model can be combined with other astrophysical models explaining the CR spectrum at different wavelengths. In particular, we provide a possible model, in which the whole CR spectrum from around GeV to 100 EeV can be explained with four main sources, including SNRs (Gal pop1), microquasars (MQs, Gal pop2), clusters of galaxies (CGs, ExGal pop1) and radio galaxies (RQs, ExGal pop2), as shown in Fig 4. Note that the main purpose here is to demonstrate that the reacceleration model works not only in the microquasars but also in the large-scale jet-cocoon system in radio galaxies.
@@
For CRs from Galactic SNRs, we calculate the observed energy spectrum of each element with DRAGON code, and the normalization is adjusted to match the AMS-02 data. We adopt a broken-power-law injection spectrum of CRs from SNRs to fit the data. See the Appendix for the details. The escaped CR spectrum typically has an exponential cutoff beyond the maximum acceleration energy. We set the maximum rigidity from Galactic SNRs to $\mathcal{R}_{\rm max}=0.4\rm~PV$.
@@
Figure 4: An overall picture of the CR spectrum from Galactic to extragalactic components. Gal pop1: CRs from SNRs Reynolds (2008); Blasi (2013) (dotted curve); Gal pop2: CRs from MQs (dashed curve); ExGal pop1: CRs from extragalactic sources, e.g., CGs Murase et al. (2008) (dotted-dashed curve); ExGal pop2: UHECRs from RGs Kimura et al. (2018a) (long dashed curve). The data points are the same as those in Fig. 3, with the AMS data taken from Refs. Aguilar et al. (2018a, b, 2015, 2021, 2020), and the mean mass distribution measured by Auger is taken from Ref. Yushkov (2020).
@@
For CRs in the knee energy range, we use the CR spectrum that comes directly from Monte Carlo simulations, as in Fig. 3. At energies above around 100 PeV to EeV, CRs from extragalactic sources begin to dominate. At this energy range, the second knee appears, where the spectrum is further steepened from $\gamma\sim 3$ to $\gamma\sim 3.3$, because of the limitation of the maximum energy of the second Galactic component based on the rigidity-dependent modeling. The physical origin of the CRs beyond the second knee up to the ankle has been proposed by various models, and it has been shown that an additional component so called the B-component is required Hillas (2005); Gaisser (2006). This component may be responsible for the proton subankle (hardening) feature around 100 PeV observed by KASCADE-Grande Apel et al. (2013) , and the proposed models include the cluster of galaxies Murase et al. (2008, 2013); Simeon et al. (2025), Galactic wind Thoudam et al. (2016); Mukhopadhyay et al. (2023), and Wolf-Rayet star explosions Thoudam et al. (2016). In addition, superknee CR accelerators such as Galactic neutron star merger remnants Kimura et al. (2018b) and interacting supernovae Ptuskin and Zirakashvili (2003, 2005) may explain another hardening feature around $10^{16.22}$ eV reported by TALE Abbasi et al. (2018).
@@
For the subankle proton component, we consider a cluster of galaxies for the demonstrative purposes. We consider that the injected CRs have a composition similar to the solar composition, but enhanced by a factor of $(A/Z)^{2}$; see the last column in Tabel 2 about $f_{\odot}^{\rm enhance}$. Note that our results are unaffected by the composition assumption unless the CRs are dominated by heavy nuclei, and such nuclei can be largely disintegrated inside the clusters Kotera et al. (2009); Fang and Murase (2018). The source spectral index above 10-100 PeV is $s_{\rm cr}=2.4$ and the maximum acceleration rigidity $\mathcal{R}_{\rm max}=3.7\times 10^{17}\rm~eV$, following Ref. Murase et al. (2008). We propagate CRs from an ensemble of extragalactic sources that are uniformly distributed with the characteristic distance between sources $d_{s}=32.5\rm~Mpc$. The number density of galaxy clusters is around $10^{-5}\rm~Mpc^{-3}$, which leads to $d_{s}\sim 46$ Mpc. The observed flux is obtained with Eq. 6 in Ref. González et al. (2021). We assumed a uniform extragalactic magnetic field $B_{\rm rms}=1\rm~nG$ and coherence length $L_{\rm coh}=1\rm~Mpc$.
@@
UHECRs above EeV energies are produced from radio galaxies via the shear acceleration mechanism Caprioli (2015); Kimura et al. (2018a); Mbarek and Caprioli (2019); Mbarek et al. (2023); Seo et al. (2024); Wang et al. (2024); Mbarek et al. (2025). Here, we adopt the predicted spectrum and composition model of UHECRs in KMZ18, where CRs with rigidity $\sim 1$ TV in radio galaxies are reaccelerated by shear in kpc-scale jets. Since radio galaxies are more evolved than our Galaxy, KMZ18 considers that the metal composition is enhanced by a ratio of 3, compared to our Galaxy. Here, we further enhance the iron composition ratio by a factor of 3 to improve the fit to the latest UHECR data. The further enhancement of the iron composition ratio in radio galaxies is reasonable because elliptical galaxies have lower star-formation rates than our Galaxy. In radio galaxies, core-collapse supernovae do not frequently occur because of their low star-formation rates, while type Ia supernovae occur because explosions of white dwarfs require significant time lag after star-formation activities. Then, type Ia supernovae inject iron into the ISM, enhancing the iron composition ratio. In addition, we consider a larger value of magnetic field strength in the cocoon with $B_{\rm coc}=4.5\rm~\mu G$, where the peak energy of the injection spectrum shown in Fig. 3 in KMZ18 is increased to $\mathcal{R}_{\rm max}\approx 2.4\rm~EV$ .
@@
We found that the overall CR spectrum and composition could be explained by the above four populations. This demonstrates that the reacceleration of Galactic CRs plays a key role in reshaping the mass distribution of accelerated CRs at different energy ranges. However, one should keep in mind that uncertainty is currently large.